# Self-assembled filament layers in drying sessile droplets: from morphology to electrical conductivity

### v1

@@ III.2.2 Percolation of filament deposits
Percolation theory describes the emergence of long-range connectivity in disordered systems and underpins transport in polymeric, colloidal, and nanowire networks [Kirkpatrick1973]. We investigate a two-dimensional network of filaments confined within a circular domain of radius $a$, where chains of length $L$ (bead diameter $d_{0}$) are deposited. Two beads are defined as connected if their center-to-center distance is less than $d_{c}=1.12d_{0}$.
We investigate filament percolation in layers deposited from evaporating droplets. To evaluate percolation, we define source and sink regions as two spherical-cap sectors located on opposite sides of the circular simulation domain, with a distance of $0.8a$. To evaluate percolation independent of hydrodynamic alignment and curvature, we generated an isotropic control system via Monte Carlo (MC) insertion of straight filaments into a periodic rectangular box matching the caps’ separation and base diameter.
Figure 7: The distance between source and sink is choosen to be $0.8a$. (a) Percolation threshold versus filament length for fully flexible and semi-stiff filaments under reaction-limited ($K=10$) and diffusion-limited ($K\approx 0$) evaporation, compared with Monte Carlo data. We observe for stiffer filaments and more uniform evaporation a decrease in the percolation thresholds. (b) Conductivity $\varkappa$ vs. area fraction $c$ for fully-flexible filaments in both evaporation regimes: $K=10$ (left) and $K\approx 0$ (right). We observe a larger critical exponent for more uniform evaporation.
@@
The percolation threshold $n_{c}$ is the critical filament number density at which a spanning cluster first appears. It is obtained as the inflection point of a sigmoid fit to the percolation probability $P(n)$ as a function of filament number density $n$ [Pawowska2013]:
@@
Here, $w$ denotes the transition width. For finite-width filaments, the critical number density $n_{c}$ decreases with increasing aspect ratio, in agreement with excluded volume theory [Balberg1984, Broman2021, Singh2025]:
$n_{c}=C\,L^{-p},$ (53)
where $C$ and $p$ are fitting parameters. In the theoretical limits, $p=2$ for very high-aspect-ratio filaments and $p=1$ for nearly isotropic particles [Balberg1984]. Since our system lies in an intermediate regime ($L/W\approx 15$), $p$ is not fixed. Here, $W$ is the filament width. The proportionality also accounts for finite-size effects and deviations from ideal scaling [Mietta2014, Li2009].
@@
Fig. 7a shows the percolation thresholds as a function of filament length of a source-sink distance of $0.8a$. Results are shown for both fully flexible and semi-stiff filaments under reaction-limited and diffusion-limited evaporation. The results are fitted by Eq. 53. These are compared with reference Monte Carlo simulations of random placed straight filaments. The corresponding fit parameters are listed in Tab. 1. For filament lengths exceeding 10, the semi-stiff case with $K=10$ yields percolation thresholds comparable to the Monte Carlo reference, as the resulting deposits consist of relatively straight, uniformly distributed filaments similar to those in the random placement model. In contrast, all other cases exhibit higher thresholds, since fully flexible filaments coil and thus have a reduced effective length. The resulting coffee-ring effect leads to lower conductivity, as it reflects only a weak transition state between uniform evaporation and a pronounced coffee-ring pattern.
Condition $\boldsymbol{C\pm\delta C}$ $\boldsymbol{p\pm\delta p}$ $\mathcal{R}^{2}$
$K\approx 0$ (fully-flex) $3.74\pm 1.28$ $1.80\pm 0.14$ 0.997
$K\approx 0$ (semi-stiff) $1.01\pm 0.21$ $1.41\pm 0.09$ 0.994
$K=10$ (fully-flex) $3.57\pm 0.70$ $1.87\pm 0.08$ 0.998
$K=10$ (semi-stiff) $7.47\pm 2.48$ $2.32\pm 0.14$ 0.998
MC – stiff $2.16\pm 0.14$ $1.82\pm 0.03$ 1.000
Table 1: Fitted parameters $C$ and $p$ (with standard errors) for Eq. 53, modelling the percolation threshold as a function of filament length. Stiffer filaments and uniform evaporation yield lower percolation thresholds. The model remains robust across flexible filaments and varying evaporation regimes, as shown by the high coefficient of determination ($\mathcal{R}^{2}\approx 1$).
Figure 8: Snapshot of deposition pattern with semi-stiff filaments at length $L=31$ under diffusion-limited evaporation ($K\approx 0$), illustrating the vapor-shielding effect between neighbouring droplets. (a) Deposition pattern of a single isolated droplet on a substrate (open circles) compared to a droplet in the presence of a neighbor (filled circles) for different area fractions $c=0.156$ (left) and $c=0.219$ (right). The neighboring droplet is located directly below at a center-to-center distance $b$ relative to the base radius $a$, with a ratio of $a/b=0.002$. (b) Smoothed normalized conductivity $\varkappa_{\xi}/\varkappa$ (squares) and area fraction $c_{\xi}/c$ (circles) averaged over the left and right half-space of the circular domain, plotted versus the azimuthal angle from the droplet center. Conductivity is evaluated perpendicular to this angle and averaged over eight independent initial configurations per area fraction (c=0.156, 0.219, 0.282, 0.345c=0.156,\,0.219,\,0.282,\,0.345). $\xi=180^{\circ}$ corresponds to the droplet–droplet axis.
@@ III.2.3 Conductivity of the Deposit
Near the percolation threshold, the conductivity as a function of concentration is expected to follow a power-law scaling [Kirkpatrick1973]:
@@
where $c=nLW$ is the area fraction, $\alpha$ the critical conductivity exponent, and $c_{t}:=n_{c}LW$ the critical area fraction. In the thermodynamic limit, this scaling arises from the scale-invariant nature of the system at the threshold [Christensen2005]. While our finite systems do not exhibit strict fractality, we expect an effective scaling behavior consistent with percolation theory.
We treat the deposited filaments as a resistor network and compute the effective conductivity by applying Kirchhoff’s laws [Kirkpatrick1973], normalizing by the source–sink distance $0.8a$. To mitigate finite-size bias, we define the background level $\varkappa_{\mathrm{bg}}$ as the conductivity observed where the percolation probability reaches $50\%$. The effective conductivity is then obtained by subtracting this baseline: $\varkappa=\varkappa_{\mathrm{raw}}-\varkappa_{\mathrm{bg}}$.
@@
Given the limited dynamic range, we adopt the power-law ansatz motivated by the percolation theory described above [Clauset2009, Sahimi2023] and extract only the exponent. Fig. 7b reports the fully flexible case under reaction- and diffusion-limited evaporation, fitted to the background-corrected data to extract the shared conductivity exponent $\alpha$ near $c_{t}$ (shown in Tab. 2). The prefactor exhibits a tentative increase with length, although large uncertainties limit the reliability of this trend, while the scaling exponent remains consistent across varying filament lengths. Those exponents deviate from the universal value depending on evaporation regime and filament stiffness, with the highest values observed for uniform evaporation in the reaction-limited regime and for stiffer filaments. A reference Monte Carlo simulation in the same domain yields $\alpha=1.26\pm 0.01$, in agreement with the universal two-dimensional conductivity exponent for random percolation ($\alpha\approx 1.3$[Berg2024] and with continuum stick-network simulations reporting $\alpha=1.24\pm 0.03$ [Balberg1983] and $\alpha=1.280\pm 0.014$ [Li2010].
@@
III.2.4 Two closely adjacent droplets
In the diffusion-limited regime ($K\approx 0$), the droplet-evaporation flux is highly sensitive to neighboring droplets, nearby obstacles, and external aerodynamic conditions [Zhang2025, Wang2022]. The evaporation profiles for interacting droplets are described by Eq. 22. The CGLB method is well-suited to efficiently resolve the fluid dynamics in these coupled systems, where analytical solutions of the evaporation profile are known. Fig. 8 illustrates how a neighboring droplet modifies the deposition pattern. In Fig. LABEL:fig:ad12, filament deposits from an isolated droplet (open circles) are compared with those from a droplet adjacent below (filled circles). For the isolated droplet, deposition is on average isotropic, reflecting symmetric capillary flows. With an adjacent droplet, vapor shielding reduces evaporation on the facing side, leading to fewer deposited filaments and a shift of deposits toward the outer boundary. Fig. LABEL:fig:ad3 presents the smoothed normalized conductivity and density (circular moving average, window size 6) as functions of the azimuthal angle, where $180^{\circ}$ denotes the direction toward the neighboring droplet. Conductivity is evaluated only within the half-space corresponding to each angle to resolve anisotropy; for example, at $180^{\circ}$, only the half-space facing the neighboring droplet is included. Consequently, both conductivity and density reach minima at $180^{\circ}$, dropping by approximately $50\%$ and $40\%$ respectively relative to the unshielded side at $0^{\circ}$. This anisotropy is tunable via inter-droplet spacing: reducing the separation intensifies vapor shielding, thereby amplifying the contrast between facing and outer boundaries. We propose using this anisotropy to experimentally identify a suitable evaporation regime.
@@ IV Conclusion
Our method was validated by simulating pure (filament-free) droplet evaporation under the spherical cap approximation in both reaction- and diffusion-limited evaporation regimes. The simulations reproduce the expected temporal evolution of droplet height and internal velocity fields.
Using this framework, we examined self-organization of the filament and final deposition patterns in the stick–slip contact line mode for medium contact angles ($\theta\approx 40^{\circ})$ and strong substrate friction. Reaction-limited evaporation yields uniform, spatially extended deposits, favorable for continuous conductive pathways in, e.g., flexible, transparent, or radio-frequency electronics. In contrast, diffusion-limited evaporation leads to pronounced coffee-ring formation. The onset of the coffee-ring formation coincides with reduced network connectivity. The orientation of deposited filaments varies across the droplet: tangential alignment dominates near the contact line, radial alignment emerges in the intermediate region, and orientations are nearly random at the center. Longer filaments preferentially align tangentially and yield more centralized deposits. The percolation threshold decreases with increasing filament length, stiffness, and with more spatially uniform evaporation (e.g., in the reaction-limited regime). The resulting deposit conductivity is well described by a fitted power-law dependence on filament concentration, with modest deviations from the expected universal scaling that depend on filament stiffness and the evaporation regime. These results are compared with Monte Carlo simulations of randomly oriented filament systems. In diffusion-limited systems of closely spaced droplets, the facing sides exhibit reduced conductivity and a more uniform but lower-density deposition, resulting from suppressed radial flow due to reduced evaporation caused by elevated vapor concentrations between the droplets. We propose this configuration for the experimental validation of the evaporation regime.
A recent study reported a link between morphology scaling and conductivity scaling [Berg2024]. Future work should test this relationship for different evaporation-driven deposits and further investigate critical behavior [Balberg1983]. Junction resistance, determined by the number and quality of inter-filament contacts, should also be considered as it can substantially affect the effective conductivity [DeStefano2019]. Beyond sessile droplets, our framework can be applied to multilayer deposits and line-shaped droplet geometries that are central to printing and coating technologies.

### v2

@@ III.2.2 Percolation of filament deposits
Percolation theory describes the emergence of long-range connectivity in disordered systems and underpins transport in polymeric, colloidal, and nanowire networks [31]. We investigate a two-dimensional network of filaments confined within a circular domain of radius $a$, where chains of length $L$ (bead diameter $d_{0}$) are deposited. Two beads are defined as connected if their center-to-center distance is less than $d_{c}=1.12d_{0}$.
We investigate filament percolation in layers deposited from evaporating droplets. To evaluate percolation, we define source and sink regions as two spherical-cap sectors located on opposite sides of the circular simulation domain, with a distance of $0.8a$. To compare to percolation in the absence of hydrodynamic alignment and curvature, we generated an isotropic control system via Monte Carlo (MC) insertion of straight filaments into a periodic rectangular box matching the caps’ separation and base diameter.
@@
The percolation threshold $n_{c}$ is the critical filament number density at which a spanning cluster first appears. It is obtained as the inflection point of a sigmoid fit to the percolation probability $P(n)$ as a function of filament number density $n$ [54]:
@@
Here, $w$ denotes the transition width. For filaments of finite aspect ratio, the critical number density $n_{c}$ decreases with increasing aspect ratio, in agreement with excluded volume theory [4, 10, 61]:
$n_{c}=C\,L^{-p}{\color[rgb]{0,0,0}+C_{0}},$ (52)
where $C$ is a fitting parameter. $p=2$ is a theoretical value for the case of high aspect-ratio filaments [4, 47, 40] and $C_{0}$ accounts for finite-size effects outside of the transient regime for high aspect ratios.
Figure 7: The distance between source and sink is chosen to be $0.8a$. (a) Percolation threshold versus filament length for fully flexible and semi-stiff filaments under reaction-limited ($K=10$) and diffusion-limited ($K\approx 0$) evaporation, compared with Monte Carlo data. We observe for stiffer filaments and more uniform evaporation a decrease in the percolation thresholds. (b) Conductivity $\varkappa$ vs. area fraction $c$ for fully-flexible filaments in both evaporation regimes: $K=10$ (left) and $K\approx 0$ (right). We observe a larger critical exponent for more uniform evaporation.
@@
Fig. 7a shows the percolation thresholds as a function of filament length for a source–sink distance of $0.8a$. For short filaments, percolation events become increasingly rare and could therefore not be sampled with sufficient statistical accuracy. Results for fully flexible and semi-stiff filaments under reaction- and diffusion-limited evaporation are fitted using Eq. 52, with the finite-size bias $C_{0}=1\times 10^{-4}$ determined from the Monte Carlo data and fixed for the remaining parameters. The fit parameters are listed in Tab. 1. For filament lengths exceeding 10, the semi-stiff case with $K=10$ yields percolation thresholds comparable to the Monte Carlo reference, as the resulting deposits consist of relatively straight, uniformly distributed filaments similar to those in the random placement model. In contrast, all other cases exhibit higher thresholds, either due to the reduced effective length of coiled, flexible filaments or due to radially inhomogeneous deposits that lie between a homogeneous film and a fully developed coffee-ring structure. In such morphologies, network connectivity is reduced because the central region becomes depleted before the peripheral accumulation is sufficiently dense to establish percolation, leading to a dip in conductivity.
Condition $\boldsymbol{C\pm\delta C}$
$K\approx 0$ (fully-flex) $5.09\pm 0.18$
$K\approx 0$ (semi-stiff) $3.93\pm 0.07$
$K=10$ (fully-flex) $4.19\pm 0.16$
$K=10$ (semi-stiff) $3.19\pm 0.07$
MC – stiff $3.29\pm 0.03$
Table 1: Fitted values of the parameter $C$ with standard errors from Eq. 52, which models the percolation threshold as a function of filament length. The finite size bias $C_{0}=1\times 10^{-4}$ was determined for the MC data and fixed for the other parameters. The results indicate that stiffer filaments and uniform evaporation conditions yield to lower percolation thresholds.
@@ III.2.3 Conductivity of the Deposit
Near the percolation threshold, the conductivity as a function of concentration is expected to follow a power-law scaling [31]:
@@
where $c=nLW$ is the area fraction, $\alpha$ the critical conductivity exponent, and $c_{t}:=n_{c}LW$ the critical area fraction. Here, $W{\color[rgb]{0,0,0}=2d_{0}}$ is the filament width. In the thermodynamic limit, this scaling arises from the scale-invariant nature of the system at the threshold [12]. While our finite systems do not exhibit strict fractality, we expect an effective scaling behavior consistent with percolation theory.
We treat the deposited filaments as a resistor network and compute the effective conductivity by applying Kirchhoff’s laws [31], normalizing by the source–sink distance $0.8a$. To mitigate finite-size bias, we define the background level $\varkappa_{\mathrm{bg}}$ as the conductivity observed where the percolation probability reaches $50\%$ [62]. The effective conductivity is then obtained by subtracting this baseline: $\varkappa=\varkappa_{\mathrm{raw}}-\varkappa_{\mathrm{bg}}$.
@@
Given the limited dynamic range, we adopt a power-law ansatz motivated by percolation theory [13, 57] and focus on extracting the scaling exponent. Fig. 7b reports the fully flexible case under reaction- and diffusion-limited evaporation, fitted to the background-corrected data to extract the shared conductivity exponent $\alpha$ near $c_{t}$ (shown in Tab. 2). These exponents deviate from the universal Monte Carlo reference value and exhibit weak trends with evaporation regime and filament stiffness, with partially overlapping confidence intervals. The largest values occur for uniform evaporation in the reaction-limited regime, particularly for fully flexible filaments. These exponents deviate from the universal Monte Carlo value and show weak trends with evaporation regime and filament stiffness. The deviation likely arises from microscopic network effects: filament flexibility and evaporation-driven aggregation create local clustering and correlations, modifying connectivity and slightly shifting the effective conductivity exponent compared to ideal percolation.
Figure 8: Snapshot of deposition pattern with semi-stiff filaments at length $L=31$ under diffusion-limited evaporation ($K\approx 0$), illustrating the vapor-shielding effect between neighbouring droplets. (a) Deposition pattern of a single isolated droplet on a substrate (open circles) compared to a droplet in the presence of a neighbor (filled circles) for different area fractions $c=0.156$ (left) and $c=0.219$ (right). The neighboring droplet is located directly below at a center-to-center distance $b$ relative to the base radius $a$, with a ratio of $a/b=0.002$. (b) Smoothed normalized conductivity $\varkappa_{\xi}/\varkappa$ (squares) and area fraction $c_{\xi}/c$ (circles) averaged over the left and right half-space of the circular domain, plotted versus the azimuthal angle from the droplet center. Conductivity is evaluated perpendicular to this angle and averaged over eight independent initial configurations per area fraction (c=0.156, 0.219, 0.282, 0.345c=0.156,\,0.219,\,0.282,\,0.345). $\xi=180^{\circ}$ corresponds to the droplet–droplet axis. Both quantities are normalized by their respective angle-averaged values.
III.2.4 Two closely adjacent droplets
A reference Monte Carlo simulation in the same domain yields $\alpha=1.26\pm 0.01$, in agreement with the universal two-dimensional conductivity exponent for random percolation ($\alpha\approx 1.3$[8] and with continuum stick-network simulations reporting $\alpha=1.24\pm 0.03$ [5] and $\alpha=1.280\pm 0.014$ [41].
@@ III.2.4 Two closely adjacent droplets
In the diffusion-limited regime ($K\approx 0$), the droplet-evaporation flux is highly sensitive to neighboring droplets, nearby obstacles, and external aerodynamic conditions [73, 66]. The evaporation profiles for interacting droplets are described by Eq. 22. The CGLB method is well-suited to efficiently resolve the fluid dynamics in these coupled systems, where analytical solutions of the evaporation profile are known. Fig. 8 illustrates how a neighboring droplet modifies the deposition pattern. In Fig. LABEL:fig:ad12, filament deposits from an isolated droplet (open circles) are compared with those from a droplet adjacent below (filled circles). For the isolated droplet, deposition is on average isotropic, reflecting symmetric capillary flows. With an adjacent droplet, vapor shielding reduces evaporation on the facing side, leading to fewer deposited filaments and a shift of deposits toward the non-shielded side. Fig. LABEL:fig:ad3 presents the smoothed normalized conductivity and area fraction (circular moving average, window size 6), where both quantities are normalized by their respective angle-averaged values over the full circular domain to highlight relative angular variations, as functions of the azimuthal angle. Where $180^{\circ}$ denotes the direction toward the neighboring droplet. Conductivity is evaluated only within the half-space corresponding to each angle to resolve anisotropy; for example, at $180^{\circ}$, only the half-space facing the neighboring droplet is included. This definition neglects percolating pathways that cross between half-spaces and therefore slightly underestimates the absolute conductivity; however, it is justified here as a measure of the angle-dependent average value. Consequently, both conductivity and area fraction reach minima at $180^{\circ}$, dropping by approximately $50\%$ and $40\%$ respectively relative to the unshielded side at $0^{\circ}$. This anisotropy is tunable via inter-droplet spacing: reducing the separation intensifies vapor shielding, thereby amplifying the contrast between facing and outer boundaries. We propose that the angular anisotropy in conductivity and filament area fraction can serve as an experimental indicator of the evaporation regime, as its magnitude directly reflects the strength of vapor shielding and depends on inter-droplet spacing.
@@ IV Conclusion
Our method was validated by simulating pure (filament-free) droplet evaporation under the spherical cap approximation in both reaction- and diffusion-limited evaporation regimes. The simulations reproduce the expected temporal evolution of droplet height and internal velocity fields. Using this framework, we examined self-organization of the filament and final deposition patterns for dilute filament suspensions in the stick–slip contact line mode for medium contact angles ($\theta\approx 40^{\circ})$, strong substrate friction, and limited filament-droplet size ratios. Reaction-limited evaporation yields uniform, spatially extended deposits, favorable for continuous conductive pathways in, e.g., flexible, transparent, or radio-frequency electronics. In contrast, diffusion-limited evaporation leads to pronounced coffee-ring formation. The onset of the coffee-ring formation coincides with reduced network connectivity. The orientation of deposited filaments varies across the droplet: tangential alignment dominates near the contact line, radial alignment emerges in the intermediate region, and orientations are nearly random at the center. Longer filaments preferentially align tangentially and yield more centralized deposits. The percolation threshold decreases with increasing filament length, stiffness, and with more spatially uniform evaporation (e.g., in the reaction-limited regime) within the concentration range considered here. The resulting deposit conductivity is well described by a fitted power-law dependence on filament concentration, with modest deviations from the expected universal scaling that depend on filament stiffness and the evaporation regime. These results are compared with Monte Carlo simulations of randomly oriented filament systems. In diffusion-limited systems of closely spaced droplets, the facing sides exhibit reduced conductivity and a more uniform but lower-density deposition, resulting from suppressed radial flow due to reduced evaporation caused by elevated vapor concentrations between the droplets. We propose this configuration for the experimental validation of the evaporation regime.
A recent study reported a link between morphology scaling and conductivity scaling [8]. Future work should test this relationship for different evaporation-driven deposits and further investigate critical behavior [5]. Junction resistance, determined by the number and quality of inter-filament contacts, should also be considered as it can substantially affect the effective conductivity [18]. Additionally, it is important to investigate how filament entanglement and variations in the receding contact angle affect the deposited morphology. Higher-resolution analysis would also enable a more detailed study of filament alignment near the contact line. Beyond sessile droplets, our framework can be applied to multilayer deposits and line-shaped droplet geometries that are central to printing and coating technologies.